OpenStudy (annie12m):

Amber measured the length of her yard by walking from one end of the yard to the other. She counted 44 steps. Assuming that each of her steps was 1'2 m in length, Amber said the length of her yard was 44 * 1/2 = 22 m. The actual length of her yard is 19 m. What was the percent error in Amber’s measurement? Round to the nearest tenth of a percent, if necessary. A. 13.6% B. 15.8% C. 76% D. 131.6%

OpenStudy (annie12m):

| 19 - 22 |= 3 3 / 19 = 0.15789473 0.15789473 * 100 = 15.789473 15.789473 = 16% Is 16% correct?

OpenStudy (annie12m):

@jayjaysimmer @pooja195 @purple_pink @RhondaSommer @silveralchemist09 @truwhovian11

OpenStudy (annie12m):

no 15.8 % i mean

OpenStudy (anonymous):

thats correct

OpenStudy (annie12m):

ok ty

OpenStudy (anonymous):

youre welcome